OpenStudy (anonymous):

if x+(2/x) = 1; then find (x^2+x+2)/(x^2(1-x))

OpenStudy (anonymous):

we have \(\Large x+\frac{2}{x}=1\)\(\Large \Leftrightarrow x^2-x+2=0\\\Large \Leftrightarrow x^2=x-2 \Rightarrow x^3=x^2-2x\) Thus, \(\Large \frac{x^2+x+2}{x^2(1-x)}=\frac{x^2+x+2}{x^2-x^3}=\frac{(x-2)+x+2}{x^2-(x^2-2x)}=\frac{2x}{2x}=1\)

OpenStudy (anonymous):

thank you soo much :) do you have any online resource to practise questions like these? how do you do these type of questions?

OpenStudy (anonymous):

I do all type which I was studied

OpenStudy (anonymous):

okey :)